OpenStudy (anonymous):

Solve 4(x + 1) = 4x + 10. 0 6 All real numbers No solution

OpenStudy (anonymous):

and why

OpenStudy (mathstudent55):

First, use the distributive property on the left side. Can you do that?

OpenStudy (anonymous):

im lost

OpenStudy (mathstudent55):

Ok. Let's work on the left side. Since we can't add x and 1 because they are not like terms, we can simplify the left side using the distributive property.

OpenStudy (anonymous):

okay go on

OpenStudy (mathstudent55):

Here is the distributive property: \(a(b + c) = ab + ac\) Here are two examples: \(2(3 + 5) = 2 \times 3 + 2 \times 5 = 6 + 10 = 16\) \(5(x + 4) = 5 \times x + 5 \times 4 = 5x + 20\)

OpenStudy (mathstudent55):

Look especially at the second example. Now we apply that to the left side of your equation.

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

\( 4(x + 1) = 4x + 10\) We use the distributive proerty on the left side to get: \(4x + 4 = 4x + 10\) Do you follow so far?

OpenStudy (anonymous):

like this

OpenStudy (anonymous):

4x+4 = 4x + 10 4 =10 , which is a false inequality is no solution

OpenStudy (mathstudent55):

Correct. The next step is to subtract 4x from both sides which gives you \(4 = 10\) Since this is a false statement, that means no solution. You got it!

OpenStudy (anonymous):

thankss

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

can u help me on more

OpenStudy (mathstudent55):

I can help you with one more. Pls start a new post. I'll look for it.

OpenStudy (anonymous):

okay thanks